Consider a single activity “Precast Wall Panels” of a certain project. Assume that the quantity of this activity is 2000 ft2 of 8’ X 8’ X 4” wall panels. Assume that crew C-11 will be used for this activity and that this crew is composed of one foreman, 6 workers, one equipment operator, and one 150-ton crane. Use the following data to estimate the unit price and total price of this activity and the time needed to finish the activity: - The daily output for crew C-11 is 320 ft2 . - Direct hourly rates for crew C-11 are as follows: - Foreman: $45 per hour. - Worker: $40 per hour. - Equipment operator: $60 per hour. - 150-ton crane: $235 per hour. - Material cost is $20/ft2 . - Total direct cost for this project: $1,000,000. - Project overheads: $100,000. - General overheads to be distributed over all company projects: $2,000,000. - Total direct cost for all projects: $10,000,000. - Assume 10% profit. - Assume 8 work hours per day..
Answer #1

Hi,

Solution as follows:

STEP 1: Calculation of percentage of project overheads:

Given Data:

  • Total Direct costs for this project = $1,000,000
  • Project overheads = $100,000

STEP 2: Calculation of percentage of general overheads:

Given Data:

  • General overheads = $2,000,000
  • Total Direct costs for all projects = $10,000,000

STEP 3: Calculation of direct cost of crew for one hour:

Given Data:

  • Crew C-11 Size = 1 foreman + 6 workers + 1 operator + 1 crane
  • Rate of foreman = $45 per hr
  • Rate of worker = $40 per hr
  • Rate of operator = $60 per hour
  • Rate of Crane = $235 per hour

Hourly cost of crew calculated in below table:

Sl. No.

Particulars

Unit Rate

Unit

Quantity

Cost

1

foreman

$45

per hour

1

$45

2

worker

$40

per hour

6

$240

3

operator

$60

per hour

1

$60

4

Crane

$235

per hour

1

$235

Total cost per hour

$580

STEP 4: Calculation of required working hours of crew:

Given Data:

  • Total Quantity = 2000 sq.ft
  • Daily output of crew = 320 sq.ft per day
  • Working hours = 8 hours per day

STEP 5: Estimation of unit price and total price for the activity:

Given Data:

  • Total Quantity = 2000 sq.ft
  • Material cost = $20 per sq.ft
  • Profit = 10%
  • General overheads = 20% [Calculated above]
  • Project overheads = 10% [Calculated above]
  • Required time = 50 hours[Calculated above]
  • Crew rate = $580 per hour[Calculated above]

Total price and unit price calculated in below table:

Sl. No.

Particulars

Unit Rate

Unit

Quantity

Cost

1

Material Cost of Panels

$20

sq.ft

2000

$40,000

2

Labor Cost of Crew C-11

$580

per hour

50

$29,000

Total Direct Costs

A

$69,000

General overheads @ 20%

B

=

Ax20%

$13,800

Project overheads @ 10%

C

=

Ax10%

$6,900

Total Costs (Direct + Indirect)

D

=

A+B+C

$89,700

Profit @ 10% of total costs

E

=

Dx10%

$8,970

Total Price for the activity

F

=

D+E

$98,670

Total Quantity (Sq.ft)

G

2,000

Unit Price (Per sq.ft)

H=F/G

$49.34

Answers:

Total time = 50 hours i.e. 6 days 2 hours

Total price = $98,670

Unit price = $49.34 per sq.ft